Presentation is loading. Please wait.

Presentation is loading. Please wait.

Week 15 LHC progress - summary Coordination: Mike Lamont, Wolfgang Hofle, Stefano Redaelli Jorg Wenniger, Massimo Gianvannozzi, Jan Uythoven.

Similar presentations


Presentation on theme: "Week 15 LHC progress - summary Coordination: Mike Lamont, Wolfgang Hofle, Stefano Redaelli Jorg Wenniger, Massimo Gianvannozzi, Jan Uythoven."— Presentation transcript:

1 Week 15 LHC progress - summary Coordination: Mike Lamont, Wolfgang Hofle, Stefano Redaelli Jorg Wenniger, Massimo Gianvannozzi, Jan Uythoven

2

3 Week 15 System commissioning with beam – Collimation – Beam dump – Feedbacks – Beam instrumentation – Machine protection – RF – Transverse damper – Injection Machine characterization – Optics measurement and correction – Magnetic machine Operations – Injection and 450 GeV – Ramp – Debugging lots and lots and lots of stuff With probe bunches

4 Collimation Coarse settings at 450 GeV in place and tested with loss maps – Propagated into ramp – Loss maps repeated with ADT blow-up First automatic BPM alignment in the LHC! – TCTPH.4L1.B1 BPM-based alignment in 10 seconds. Machine protection tests Stefano Redaelli, Gianluca Valentino, Belen Maria Salvachua Ferrando and team

5 RF Progress Capture etc. over Easter RF clocks adjusted for bunch one position All cavities phased using beam based measurement, individual cavity feedback loops adjusted Fine adjustment of phase shifters for cavity sum requires access to UX45 with RF on Philippe Baudrenghien, Helga Timko

6 BI - BPM BPM phasing – Some issues with FESA class (BPMLHC) for automatic phase-in Individual BPM - issues to be tackled Spikes in capture mode – Capture data shows intermittent spikes on many channels. Can also be observed in raw capture data in FESA class, where unphysical values are reported (work around – turn-off orbit mode) Rhodri Thibaut Eva …

7 Instrumentation (1) Wire-scanner

8 Instrumentation (2) Longitudinal Density Monitor (LDM) probe bunch and de-bunched beam

9 Instrumentation (3) Federico Roncarolo, Enrico Bravin, Georges Trad et al Synchrotron light monitors

10 BLMs: loss maps with ADT Bernd Dehning, Christos Zamantzas and team Looking good

11 Feedbacks Orbit – Issues with response matrix calculation – still be resolved – OK for ramp (single optics) with some loving care and attention Tune – High sensitivity giving reasonable signal and used successfully in first part of ramp

12 ADT Progress ADT blow-up with noise commissioned – new VME hardware developed for run 2 capable to combine ultimately signals of four pick-ups Application for blow-up re-commissioned – new FESA3 classes tested and ready to use for collimation set-up – coarse adjustment for batch-by-batch blow-up done ADT Beam position module set-up for feedback loop – noticed non-conformity in cabling – essentially swap of beam 1/ beam 2 right of IP4 – requires access to tunnel, suggested procedure sent to T. Otto for approval Gerd Kotzian, Daniel Valuch

13 OMC Suffered from teething problems with BPM phasing, capture, and AC dipole AC dipole checked, MKA used and OK – Local coupling correction verified ok Beta-beat correction – generally bare machine not quite as good as in 2012 – 1 iteration beam 1, 3 iterations beam 2, <20% – Correction incorporated into ramp Rogelio Tomas and large team

14 Optics - Beta beat OMC team Beam 1, before correction: 2015 vs. 2012

15 Optics - Beta beat OMC team Beam 1, after correction: 2015 vs. 2012

16 Injection B2 – extraction checks after COLDEX plus re-steering of line MSI Aperture Measurements injection channel: – Beam1: - aperture scan was completed during ST on March 6-8th - Parallel vertical steering completed: total aperture = 9.7 sigma (-4.4 sigma and +5.3 sigma). – Beam 2: - Full aperture scan completed: H aperture = +/- 9 sigma, V aperture = +/- 7 sigma, skew = +8 /-7 sigma - Parallel vertical steering completed: total aperture ~ 10 sigma (-5 mm and +6.5 mm). No clear bottlenecks found, detailed data analysis will follow Chiara Bracco, Wolfgang Bartmann and team

17 Beam dump system Systematic run through of LBDS machine protection tests without beam – Left from cold check-out Measured the rise-time of MKD waveform for both beams – Beam injected in different buckets inside the abort gap – Dump the beam and check position on the BTVDD and BTVSE (screen in front of the TCDS). Etienne Carlier, Nicholas Magnin

18 MKD scan (Wednesday) – (2) E. Carlier, M. Fraser, N. Magnin Beam 1 OK

19 Machine protection IST tests for the LHC and injection BIS as well as for the SMP can be considered completed (M. Zerlauth and team) Ce qu’il reste à faire « offline » (avec logging) : – Vérifier les intervalles entre les trames d’énergies pour les 2 CISR (logging de l’history buffer) – Tester un événement Post Mortem et vérifier dans le PM viewer que les données sont cohérentes Ce qu’il reste à faire pour une éventuel prochain slot de 4h : – Finir les tests du SMP du SPS (encore à préciser) Ce qu’il reste à faire avec des rampes d’énergie et/ou des faisceaux : – Vérifier la cohérence et transmission sur timing de valeurs d’intensités assez élevées – Vérifier que les connections des BPF sur les BIC d’extraction ne sont pas inversées entre Beam1 et Beam2 Another slot without beam to be scheduled plus a number of dedicated tests with beam

20 Operation Q, Q’, coupling… Orbit corrected and new reference established ("First Easter Monday pilot reference") (JW) COD polarity checks, COD response measurement, BPM checks… Decay Ramp Lots and lots of debugging

21 Chromaticity at injection

22 Ramp 1

23 Ramp 2 Possibly first collisions. Thanks to Fabiola for the limoncello.

24 Ramp 3: chromaticity in the ramp 4 people in the LHC island by this stage Start ramp

25 Ramps 1B1 to tune excursion start ramp 2B2 to SBF glitch 3Dumped by BLMs at 6.5 TeV 4B2 to SBF glitch 5Quench on flat-top S81 ( C14R8)

26 BLM beam dump 6.5 TeV With the complete PM, one can actually see a loss signal that seems coherent (time and space - multiple BLMs) for the recent dump at 6.5 TeV on BPMs in R8. There are short loss bursts (UFO or vacuum?). Jorg

27 Accesses RF – 2 hours, UX45, RF ON Transverse damper – strip line pickups located in the magnets Q7, Q8, Q9 and Q10 at both left and right sides of IP4. ATLAS – 4 hours this week AC dipole – Nicolas Magnin LBDS (and MKI after scan) – Configuration update following scans Vacuum IR7 – Work in progress – to finish BLM Will give access this morning in shadow of quench recovery

28 Conclusions Long way to go but machine looks in good shape so far Wrapping up work with probes at 450 GeV – Outstanding: MP, injection, LBDS checks, ADT – Possible first look at nominal bunch, squeeze later this week Big, big thanks to all the teams who worked holidays, weekends, nights, evenings and even days


Download ppt "Week 15 LHC progress - summary Coordination: Mike Lamont, Wolfgang Hofle, Stefano Redaelli Jorg Wenniger, Massimo Gianvannozzi, Jan Uythoven."

Similar presentations


Ads by Google